# A Sampling-based Framework for Hypothesis Testing on Large Attributed Graphs [Scalable Data Science]

### v1

@@ Abstract.
Hypothesis testing is a statistical method used to draw conclusions about populations from sample data, typically represented in tables. With the prevalence of graph representations in real-life applications, hypothesis testing in graphs is gaining importance. In this work, we formalize node, edge, and path hypotheses in attributed graphs. We develop a sampling-based hypothesis testing framework, which can accommodate existing hypothesis-agnostic graph sampling methods. To achieve accurate and efficient sampling, we then propose a Path-Hypothesis-Aware SamplEr, PHASE, an $m$-dimensional random walk that accounts for the paths specified in a hypothesis. We further optimize its time efficiency and propose $\text{PHASE}_{\text{opt}}$. Experiments on real datasets demonstrate the ability of our framework to leverage common graph sampling methods for hypothesis testing, and the superiority of hypothesis-aware sampling in terms of accuracy and time efficiency.
@@ 1. Introduction
Hypothesis testing finds widespread application in various domains such as marketing, healthcare, and social science (Newey and McFadden, 1994). Hypotheses are usually tested on representative samples since it is impractical or even impossible to extract data from entire populatations due to size, accessibility or cost. For instance, snowball sampling has been proven effective in accessing a hidden population like drug abusers through a chain-referral mechanism (Waters, 2015; Goodman, 1961). In this paper, we study hypothesis testing on graphs.
Graphs can be used to represent many real-world applications such as social, bibliographic and transportation networks where entities are nodes and edges reflect relationships between entities. Specifically, we work on Attributed Graphs that are graphs with multiple types of nodes and edges, both of which are labeled with attributes containing valuable information for graph analytics (Croft et al., 2011; Omidvar-Tehrani et al., 2019). In a bibliographic network like DBLP (see Figure 1a and Figure 1b for an example and an attributed path of length one, respectively), hypotheses expressed on nodes, edges, and paths unveil interesting aspects of collaboration behavior and research trends based on citation evolution over time. Specifically, a node hypothesis can be “the average citation of conference papers is greater than average”, and a path hypothesis can be “the minimum number of citations of conference papers written by Microsoft researchers is greater than average”. Similarly, on social platforms like Yelp, hypothesis testing can be used to check how different users review a product and study its adoption.
@@
Objectives and Challenges. Given an input attributed graph $\mathcal{G}$, our aim is to verify a hypothesis $H$ and return true or false. We hence seek to achieve two objectives: O1 - hypothesis testing should be as accurate as possible; O2 - execution time should be as small as possible. To the best of our knowledge, there is no existing literature addressing hypotheses on attributed graphs. A straightforward approach is to obtain a sampled graph $\mathcal{S}$ from $\mathcal{G}$ using hypothesis-agnostic graph sampling methods (Maiya and Berger-Wolf, 2010; Li et al., 2019; Adamic et al., 2001). However, the challenge is that depending on the sampling method and on the budget $B$, i.e., the size of $\mathcal{S}$, the hypothesis estimator, which computes aggregated values of nodes, edges, or paths specified in the hypothesis may return different values on $\mathcal{S}$ and $\mathcal{G}$. This raises the accuracy challenge behind O1. Additionally, as $B$ increases, the accuracy of the resulting hypotheses increases. However, the speed at which the accuracy converges highly depends on the sampling. This raises the efficiency challenge behind O2.
@@
Contributions. We classify hypotheses on attributed graphs into three types: node, edge, and path hypotheses. We develop a sampling-based hypothesis testing framework, which accommodates common hypothesis-agnostic sampling methods, such as random node sampler (Stumpf et al., 2005), simple random walk (Gjoka et al., 2010), and non-backtracking random walk (Lee et al., 2012).
@@
The lack of awareness of the input hypothesis in existing sampling methods may slow down the convergence speed of accuracy. Indeed, the sampled graph $\mathcal{S}$ should preserve as many relevant paths according to $H$ from $\mathcal{G}$ as possible for a path hypothesis. Therefore, to address objectives O1 and O2, we design PHASE, a Path-Hypothesis-Aware SamplEr. PHASE is aware of $H$ and preserves its nodes, edges, or paths. Consequently, the resulting $\mathcal{S}$ is more likely to provide relevant information for testing the hypothesis. PHASE employs $m\geq 1$ dependent random walks with two weight functions to ensure path-hypothesis-awareness. One function biases the seed selection toward the first node in the path hypothesis. The other biases the transition probability towards the nodes specified in the hypothesis. We further propose $\text{PHASE}_{\text{opt}}$ to improve the time efficiency of PHASE. $\text{PHASE}_{\text{opt}}$ reduces computation overhead through random sampling of neighbors. It also adopt a non-backtracking approach to prioritize the selection of relevant nodes, edges, and paths. We demonstrate both theoretically and empirically that the hypothesis estimator on $\mathcal{S}$ converges to its corresponding value in $\mathcal{G}$ as $B$ increases.
@@
We conduct extensive experiments on three real-world datasets: MovieLens, DBLP, and Yelp. We aim to demonstrate the effectiveness of the optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE, and to compare hypothesis-agnostic and hypothesis-aware samplers in terms of test significance, accuracy, and running time.
@@
We first observe that $\text{PHASE}_{\text{opt}}$ is at least 20 times faster than PHASE. Then, by comparing $\text{PHASE}_{\text{opt}}$ with 11 state-of-the-art hypothesis-agnostic samplers, we find it achieves the best accuracy given the sampling budget across various hypothesis types and datasets. $\text{PHASE}_{\text{opt}}$ also exhibits fast convergence of hypothesis estimators and robust accuracy performance especially for hypotheses with longer paths or limited relevant nodes, edges, or paths in $\mathcal{G}$. For significance, $\text{PHASE}_{\text{opt}}$ consistently provides the most precise and reliable estimate, achieving the highest hypothesis significance among all samplers. Moreover, $\text{PHASE}_{\text{opt}}$ achieves high accuracy in the shortest amount of time, making it usable in practice.
@@ 2.1. Attributed Graphs
The DBLP Network in Figure 1a is an attributed graph containing four types of nodes and four types of edges. For example, paper nodes belong to the node type $\mathtt{paper}\in\mathcal{T}$ with five attributes, including title and year. The edge from paper to field of study (FOS) nodes is denoted by $\mathtt{WithDomain}\in\mathcal{R}$, with a weight indicating the paper’s relevance to a particular FOS. The reverse relationship $\mathtt{WithDomain}^{-1}\in\mathcal{R}$ also holds from FOS to paper nodes.
@@
An attributed path is a path where each node has a list of attributes, referred to as a modifier. Figure 1b presents an example of a length-one attributed path, “conference papers written by Microsoft researchers”, and Figure 1c is an instance of that path.
@@ 2.2. Hypotheses on Attributed Graphs
[…]
@@
[…]
@@
When $\mathcal{P}$ has length zero, a path hypothesis is reduced to a node hypothesis. An example of node hypothesis in the DBLP dataset can be “the average number of citations for conference papers is larger than 50” expressed as:
[…]
When $\mathcal{P}$ has length one, we refer it as an edge hypothesis. For instance, the edge hypothesis “the average $\mathtt{FOS\_weight}$ of conference papers on data mining is larger than 0.5” can be expressed as:
[…]
@@ 2.3. Problem Statement and Challenges
Given an attributed graph $\mathcal{G}$, our aim is to verify a node, edge, or path hypothesis $H$ and return true or false.
To address our problem, we formulate two objectives: O1 the hypothesis testing should be as accurate as possible; O2 the overall execution time should be as small as possible.
When testing a hypothesis such as “the average number of citations for conference papers is larger than 50”, a common approach includes collecting representative sample data and computing an aggregate value. In this case, the aggregate value corresponds to the average number of citations. However, directly collecting all relevant, i.e., satisfying the modifiers in the hypothesis, nodes, edges, or paths from $\mathcal{G}$ and computing an aggregate value can be time-consuming. Therefore, we adopt the commonly used approach of sampling the graph $\mathcal{G}$ into a smaller representative graph $\mathcal{S}$. This assumes a sampling budget $B$ that reflects the maximum size of $\mathcal{S}$. For the sake of simplicity, unless stated otherwise, we assume sampling an edge or a node requires the same unitary cost $1$.
A hypothesis estimator returns the aggregated value of nodes, edges, or paths in a hypothesis. The challenge of addressing O1 is that depending on the sampling method and the budget $B$, the hypothesis estimator obtained from $\mathcal{S}$ may differ from that of $\mathcal{G}$. We propose to first examine the applicability of existing sampling methods for graphs. Our observation is that the accuracy of those methods increases slowly as $B$ increases, resulting in longer times to achieve satisfactory accuracy. We conjecture that this is attributed to the lack of awareness of the underlying hypothesis during sampling. For instance, $\mathcal{S}$ should preserve as many relevant paths from $\mathcal{G}$ as possible for a path hypothesis. Hence, to address objective O2, we seek to design a sampling method that is aware of $H$ and preserves relevant nodes, edges, or paths in $H$. We further aim to demonstrate both theoretically and empirically that using our sampling, the hypothesis estimator on resulting $\mathcal{S}$ converges to its corresponding value in $\mathcal{G}$ as $B$ increases.
@@ 3. Sampling-based Hypothesis Testing
We propose a sampling-based hypothesis testing framework that supports both hypothesis-agnostic and hypothesis-aware graph samplers. After reviewing existing hypothesis-agnostic methods, we introduce our hypothesis-aware sampler, PHASE, in Section 3.2.1. It addresses objective O1 by ensuring convergence of hypothesis estimators, and objective O2 by incorporating hypothesis awareness into the sampling. Additionally, we implement two optimizations, $\text{PHASE}_{\text{opt}}$, to reduce the running time of PHASE in Section 3.2.2.
@@ 3.1. Hypothesis-Agnostic Samplers
Figure2 illustrates the sampling-based hypothesis testing framework we propose to test node, edge, and path hypotheses on attributed graphs. Based on whether the graph sampling method is hypothesis-agnostic or hypothesis-aware, there are two workflows indicated by dashed and dash-dot-dash arrows, respectively. Their distinction lies in where the hypothesis $H$ is introduced. In the hypothesis-aware workflow, $H$ is necessary during the sampling step, while in the hypothesis-agnostic workflow, $H$ is directly introduced before hypothesis testing.
Following dashed arrows from the left, the first procedure is to sample a graph $\mathcal{S}$ using a hypothesis-agnostic sampler, which requires $\mathcal{G}$ and $B$ only. The existing hypothesis-agnostic samplers can be categorised into three categories: node samplers that choose $B$ nodes from $\mathcal{G}$(Stumpf et al., 2005; Adamic et al., 2001), edge samplers that choose $B$ edges from $\mathcal{G}$(Krishnamurthy et al., 2005), and random walk based samplers that pick edges and nodes through the process of random walks(Gjoka et al., 2010; Lee et al., 2012; Leskovec and Faloutsos, 2006; Li et al., 2019). The second procedure is to conduct the hypothesis testing using $\mathcal{S}$ and $H$. Specifically, we extract all relevant nodes, edges, or paths from $\mathcal{S}$ and compute the required aggregated values for hypothesis testing. Finally, the acceptance or rejection result, p-value, and confidence interval are returned.
@@ 3.2.1. PHASE Algorithm
We introduce PHASE, our Path-Hypothesis-Aware SamplEr (Algorithm1). PHASE aims to preserve nodes, edges, or paths specified in a hypothesis $H$. It can be integrated with any random walk based sampler that maintains the path-preserving property. To ease understanding, we illustrate it with FrontierS(Ribeiro and Towsley, 2010). FrontierS picks a node from $m$ random seeds based on degree-proportional probability and performs a random walk by uniformly selecting a neighbor to proceed. It repeats the process until reaching the sampling budget $B$.
@@
1:Input: $\mathcal{G}=(\mathcal{V,E})$, $B$, $H$
@@
4:$L_{w}=\text{AssignWeight}(L,H)$triangleright Assign $w_{h}$ to $x_{1}$ nodes and $w_{l}$ to other nodes.
@@
10:    $N_{w}=\text{AssignWeight}(N,H)$triangleright Assign weights to nodes in $N$ according to the transition probability matrices.
11:    Select $u\in N$ with probability $N_{w}$
@@
PHASE takes $\mathcal{G}$, $B$, and $H$ as inputs. For a node, edge, or path hypothesis, $H$ contains $\mathcal{P}$ with lengths zero, one, or more than one, respectively. It initializes $m$ seed nodes randomly in a list $L$ (Line 1), which increases the probability of selecting relevant nodes, edges, and paths from $\mathcal{G}$. Each seed is assigned a weight in $L_{w}$ (Line 2) to guide the selection of seed nodes for random walks later. These weights are determined based on heuristics, with nodes satisfying the first node modifier on $\mathcal{P}$ (denoted by $x_{1}$) assigned a higher weight $w_{h}$ and all other nodes assigned a lower weight $w_{l}$.
Lines 4-14 describe an iterative random walk to select nodes until $B$ is reached. During each iteration, the algorithm chooses a node $v$ from $L$ with probability $L_{w}$ (Line 6). Then, it selects one of $v$’s neighbors $u$ using a weighted selection process (Lines 7-9). Specifically, in Line 9, we adjust the transition probability to bias the random walks towards nodes, edges, or paths that satisfy $\mathcal{P}$ in $H$. The transition probability matrices for node, edge, and path hypotheses are depicted in Figures 3a, 3b, and 3c, respectively.
After adding nodes $u$ and $v$ to $\mathcal{V}_{S}$ (Line 10), the algorithm updates $L$ by replacing $v$ with $u$ (Lines 11) and decreases the sampling budget by one (Lines 12). This iteration continues until $B>m$. The resulting sampled graph $\mathcal{S}$ is the induced subgraph from $\mathcal{V}_{S}$.
$x_{1}$$y$$x_{1}$$w_{h}$$w_{l}$$y$$w_{h}$$w_{l}$(a)$x_{1}$$x_{2}$$y$$x_{1}$$w_{l}$$w_{h}$$w_{l}$$x_{2}$$w_{h}$$w_{l}$$w_{l}$$y$$w_{h}$$w_{l}$$w_{l}$(b)$x_{1}$$x_{2}$$x_{3}$$y$$x_{1},x_{1}$$w_{l}$$w_{h}$$w_{l}$$w_{l}$$x_{1},x_{2}$$w_{l}$$w_{l}$$w_{h}$$w_{l}$$x_{1},x_{3}$$w_{h}$$w_{l}$$w_{l}$$w_{l}$$x_{1},y$$w_{h}$$w_{l}$$w_{l}$$w_{l}$$\vdots$(c)
Figure 3. Transition probability matrices for node (a), edge (b), and path ($l=2$) (c) hypotheses, where $w_{h}\geq w_{l}>0$ denote transition probabilities. $x_{i}$ represents nodes in $\mathcal{G}$ satisfying the $i$-th node modifier on $\mathcal{P}$, while $y$ represents nodes not satisfying any node modifier on $\mathcal{P}$. (a) and (b) involve 1st-order random walks, whereas (c) involves 2nd-order random walks because the probability of selecting the next node depends on both the current and previous nodes.

### v2

@@ Abstract.
Hypothesis testing is a statistical method used to draw conclusions about populations from sample data, typically represented in tables. With the prevalence of graph representations in real-life applications, hypothesis testing on graphs is gaining importance. In this work, we formalize node, edge, and path hypotheses on attributed graphs. We develop a sampling-based hypothesis testing framework, which can accommodate existing hypothesis-agnostic graph sampling methods. To achieve accurate and time-efficient sampling, we then propose a Path-Hypothesis-Aware SamplEr, PHASE, an $m$-dimensional random walk that accounts for the paths specified in the hypothesis. We further optimize its time efficiency and propose $\text{PHASE}_{\text{opt}}$. Experiments on three real datasets demonstrate the ability of our framework to leverage common graph sampling methods for hypothesis testing, and the superiority of hypothesis-aware sampling methods in terms of accuracy and time efficiency.
PVLDB Reference Format:
PVLDB, 17(11): 3192 - 3200, 2024.
doi:10.14778/3681954.3681993
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 17, No. 11 ISSN 2150-8097.
doi:10.14778/3681954.3681993
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/Carrieww/GraphHT.
@@ 1. Introduction
Hypothesis testing finds widespread application in various domains such as marketing, healthcare, and social science (Newey and McFadden, 1994). Hypotheses are usually tested on representative samples since it is impractical, or even impossible to extract data from entire populations due to size, accessibility or cost. For instance, snowball sampling has been proven effective in accessing a hidden population like drug abusers through a chain-referral mechanism  (Waters, 2015; Goodman, 1961). In this paper, we study hypothesis testing on graphs.
@@
Graphs can represent real-world applications such as social, bibliographic, and transportation networks where entities are nodes and edges reflect relationships between them. We focus on attributed graphs that contain multiple types of nodes and edges, each labeled with attributes that provide valuable information for graph analytics (Croft et al., 2011; Omidvar-Tehrani et al., 2019). In a bibliographic network like DBLP (see Figure 1a), hypotheses expressed on nodes, edges, and paths unveil interesting aspects of collaboration behavior and research trends based on citation patterns. A node hypothesis examines attributes of a single node type. For example, “the average citation of conference papers is greater than average” concerns the citation attribute of the conference paper nodes. Path hypotheses consider attributes along a defined path, such as a path connecting authors from Microsoft to their conference papers (see Figure 1b and 1c for the path and an path instance, respectively). An example of path hypothesis is “the minimum number of citations of conference papers written by Microsoft researchers is greater than average”.
@@
Objectives and Challenges. Given an attributed graph $\mathcal{G}$, our aim is to verify a hypothesis $H$ and return a true or false result. We aim to achieve two objectives: O1 - ensure hypothesis testing is as accurate as possible; O2 - minimize execution time. To the best of our knowledge, there is no existing literature addressing hypotheses on attributed graphs. A straightforward approach is to sample a subgraph $\mathcal{S}$ from $\mathcal{G}$ using hypothesis-agnostic sampling methods (Maiya and Berger-Wolf, 2010; Li et al., 2019; Adamic et al., 2001). Sampling is important when the full graph is not accessible or when the hypothesis targets very specific nodes, edges, or paths that are time-consuming to collect. However, hypothesis-agnostic samplers struggle to achieve both objectives. First, hypothesis-agnostic samplers may miss relevant nodes, edges, or paths, i.e., those requested by the hypothesis, especially when the sampling budget $B$, i.e., the size of $\mathcal{S}$, is small. The irrelevant nodes, edges, or paths, which must be filtered out for hypothesis testing, can compromise accuracy. This raises the accuracy challenge behind O1. We show in Section 3.2.3 that the hypothesis estimator, which computes aggregate values of nodes, edges, or paths specified in $H$, converges to the ground truth in $\mathcal{G}$ as $B$ increases, resulting in accuracy approaching one. However, the rate of convergence depends on the sampler, presenting a time efficiency challenge behind O2. By optimizing the sampler design, we aim to balance accuracy and time efficiency, enabling earlier sampling halts with accurate hypothesis testing results.
@@
Contributions. We classify hypotheses on attributed graphs into three types: node, edge, and path ones. We develop a sampling-based hypothesis testing framework, which accommodates common hypothesis-agnostic samplers, such as random node sampler (Stumpf et al., 2005), random walk (Gjoka et al., 2010), and non-backtracking random walk (Lee et al., 2012).
@@
The lack of awareness of the input hypothesis in existing sampling methods may slow down the rate of convergence of accuracy. Therefore, to address objectives O1 and O2, we design PHASE, a Path-Hypothesis-Aware SamplEr. PHASE is aware of $H$ and preserves the corresponding nodes, edges, or paths. Consequently, the resulting $\mathcal{S}$ is more likely to accurately test the hypothesis. PHASE employs $m\geq 1$ dependent random walks with two weight functions to ensure path-hypothesis-awareness. One function prioritizes the seed selection toward the first node in the path hypothesis. The other steers the transition probability towards the nodes specified in the hypothesis. We further propose $\text{PHASE}_{\text{opt}}$ to improve the time efficiency of PHASE. $\text{PHASE}_{\text{opt}}$ adopts a non-backtracking approach to avoid selecting previously visited nodes. It also reduces computation overhead by fixing the number of neighbors to examine. We show both theoretically and empirically that, for all samplers, the hypothesis estimator converges to the ground truth as $B$ increases, and our proposed samplers ensures earlier and smoother convergence of the hypothesis estimator.
@@
We conduct extensive experiments on three real-world datasets: MovieLens, DBLP, and Yelp. We aim to demonstrate the effectiveness of two optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE, and to compare the test significance, accuracy, and execution time of hypothesis-agnostic and hypothesis-aware samplers.
@@
We observe $\text{PHASE}_{\text{opt}}$ is at least 43 times faster than PHASE with less than 4% accuracy difference in DBLP. For significance, $\text{PHASE}_{\text{opt}}$ consistently delivers the most precise and reliable estimates. Compared to 11 state-of-the-art hypothesis-agnostic samplers, we find $\text{PHASE}_{\text{opt}}$ excels in accuracy when $B$ is fixed across various hypothesis types and datasets. It also demonstrates robust accuracy performance, especially for difficult hypotheses, i.e., those with longer paths or fewer relevant nodes, edges, or paths in $\mathcal{G}$. Moreover, the high accuracy achieved in the shortest amount of time makes our proposed sampler usable in practice.
@@ 2.1. Attributed Graphs
The DBLP Network in Figure 1a is an attributed graph containing four types of nodes and four types of edges. For example, paper nodes belong to the node type $\mathtt{paper}\in\mathcal{T}$ with five attributes, including title and year. Edges from paper to field of study (FOS) nodes belong to the edge type $\mathtt{WithDomain}\in\mathcal{R}$ with a weight attribute indicating the paper’s relevance to an FOS. The reverse relationship $\mathtt{WithDomain}^{-1}\in\mathcal{R}$ holds accordingly from FOS to paper nodes.
@@
An attributed path is a path where each node has some attributes, referred to as modifiers. Figure 1b presents an example of a length-one attributed path, “conference papers written by Microsoft researchers”, and Figure 1c is an instance of that path.
@@ 2.2. Hypotheses on Attributed Graphs
$H_{path}:P^{o}_{c}\left(\text{agg}(f_{\mathcal{P}}\mid M_{t_{i}},\forall t_{i}\text{ on }\mathcal{P})\right)$
where $o\in\{=,<>,>,<\}$; $c\in\mathbb{R}$ is a constant value; $P^{o}_{c}$ is the predicate in the format: equal, unequal, greater, or less than a value; $f_{\mathcal{P}}$ is any function of node and/or edge attributes on $\mathcal{P}$; $\text{agg}(f_{\mathcal{P}}\mid M_{t_{i}},\forall t_{i}\text{ on }\mathcal{P})$ is an aggregation function applying on the $f_{\mathcal{P}}$ of all paths whose nodes satisfy the corresponding modifiers.
@@
$P^{>}_{100}\left(\textit{avg}(f_{\mathcal{P}}\mid\mathtt{author}[\mathtt{MSR}],\mathtt{paper}[],\mathtt{author}[\mathtt{MSR}]\right)$
@@
When $l=0$, a path hypothesis is reduced to a node hypothesis. An example in DBLP, “the average number of citations for conference papers is larger than 50”, can be expressed as:
$P^{>}_{50}\left(\textit{avg}(\mathtt{paper}[\mathtt{citation}]\mid\mathtt{paper}[\mathtt{conference}])\right)$
When $l=1$, we refer it as an edge hypothesis. For instance, the edge hypothesis, “the average $\mathtt{FOS\_weight}$ of conference papers on data mining is larger than 0.5”, can be expressed as:
$P^{>}_{0.5}\left(\textit{avg}(\mathtt{WithDomain}[\mathtt{FOS\_Weight}]\mid\mathtt{paper}[\mathtt{C}],\mathtt{FOS}[\mathtt{DM}]\right)$
@@ 2.3. Problem Statement and Challenges
Given an attributed graph $\mathcal{G}$, our aim is to verify a node, edge, or path hypothesis $H$ and return true or false. We formulate two objectives to address our problem: O1 - ensure hypothesis testing is as accurate as possible; O2 - minimize execution time.
When testing a hypothesis like “the average number of citations for conference papers is greater than 50”, a conventional approach is to collect representative conference papers from the database, compute the average number of citations, and perform the statistical test. However, the entire graph may not be accessible, or collecting all relevant nodes, edges, or paths from $\mathcal{G}$ is time-consuming and impractical. Hence, we adopt graph sampling techniques to sample a subgraph $\mathcal{S}$ from $\mathcal{G}$, assuming a sampling budget $B$ that reflects the maximum size of $\mathcal{S}$. For simplicity, unless stated otherwise, sampling an edge or a node incurs the same unitary cost of one.
A hypothesis estimator computes an aggregate value for nodes, edges, or paths based on $H$. It directly impacts the accuracy of hypothesis testing. The challenge for hypothesis-agnostic samplers to achieve O1 is that they can miss relevant nodes, edges, or paths when $B$ is small. By examining the applicability of existing hypothesis-agnostic samplers, we observe that a large $B$ (almost the size of $\mathcal{G}$) is required to achieve an accuracy of one, which results in longer execution time and the challenge behind O2. We conjecture that this is due to a lack of awareness of the underlying hypothesis during sampling. Hence, to address two challenges, we aim to design a hypothesis-aware sampler that preserves relevant nodes, edges, or paths in $S$. We also intend to show, both theoretically and empirically, that using our sampler, the hypothesis estimator on $\mathcal{S}$ will converge earlier to its corresponding value in $\mathcal{G}$ as $B$ increases.
@@ 3. Sampling-based Hypothesis Testing
We propose a sampling-based hypothesis testing framework that supports both hypothesis-agnostic and hypothesis-aware graph samplers. After reviewing existing hypothesis-agnostic methods, we introduce our hypothesis-aware sampler, PHASE, in Section 3.2.1. It achieves O1 by incorporating hypothesis awareness into the sampling, and O2 by ensuring earlier convergence of hypothesis estimators. Additionally, we implement two optimizations in $\text{PHASE}_{\text{opt}}$ to reduce the execution time of PHASE in Section  3.2.2.
@@ 3.1. Hypothesis-Agnostic Samplers
Figure 2 illustrates our sampling-based hypothesis testing framework for testing node, edge, and path hypotheses on attributed graphs. It consists of two steps: (1) Sampling and (2) Hypothesis Testing. There are two workflows: hypothesis-agnostic, shown with dashed arrows, where $H$ is considered only in the hypothesis testing step, and hypothesis-aware, indicated by dash-dot-dash arrows, which requires $H$ during the sampling step.
In the hypothesis-agnostic workflow (dashed arrows), given $\mathcal{G}$ and $B$, a hypothesis-agnostic sampler is used in the sampling step to obtain $\mathcal{S}$. Existing hypothesis-agnostic samplers fall into three categories: node samplers that choose $B$ nodes from $\mathcal{G}$ (Stumpf et al., 2005; Adamic et al., 2001), edge samplers that choose $B$ edges from $\mathcal{G}$ (Krishnamurthy et al., 2005), and random walk based samplers that pick edges and nodes by random walks (Gjoka et al., 2010; Lee et al., 2012; Leskovec and Faloutsos, 2006; Li et al., 2019). In the hypothesis testing step, relevant nodes, edges, or paths are extracted from $\mathcal{S}$ for hypothesis testing. Finally, the acceptance result, p-value, and confidence interval are returned.
@@ 3.2.1. PHASE Algorithm
We introduce PHASE, our Path-Hypothesis-Aware SamplEr (Algorithm 1), for the sampling step. Since $\mathcal{G}$ may contain both relevant and irrelevant nodes, edges, or paths for a given $H$, PHASE aims to preferentially sample those specified in $H$. This strategy resembles stratified sampling (Cochran, 1977; Neyman, 1992), where the target population is sampled at higher rates without bias. PHASE can be integrated with any random walk based sampler. To clarify, we describe it using Frontier Sampler (FrontierS(Ribeiro and Towsley, 2010), which picks a node from $m$ random seeds based on degree-proportional probability and performs a random walk by uniformly selecting a neighboring node. This process repeats until $B$ is reached.
@@
1:Input: $\mathcal{G}=(\mathcal{V,E})$, $B$, $H$, $Q$
@@
4:$L_{w}=\text{AssignWeight}(L,H)$triangleright Assign $w_{h}$ to $x_{1}$ nodes and $w_{l}$ to others.
@@
10:    $N_{w}=\text{AssignWeight}(N,H,{Q})$triangleright Assign weights according to $Q$.
11:    Select $u\in N$ with the normalized $N_{w}$
@@
PHASE takes $\mathcal{G}$, $B$, $H$, $Q$ as inputs. For a node, edge, or path hypothesis, $H$ contains $\mathcal{P}$ with lengths zero, one, or more than one, respectively. $m$ seed nodes in $\mathcal{G}$ are randomly initialized and stored in a list $L$ (Line 1). This line increases the chance of picking relevant nodes, edges, and paths and ensures they are not locally clustered, preventing locality bias. Each seed is assigned a weight in $L_{w}$ (Line 2) to guide the selection of seed nodes for random walks later. The weights are determined based on heuristics: nodes satisfying the first node modifier on $\mathcal{P}$ (denoted by $x_{1}$) receive a higher weight $w_{h}$, while others receive a lower weight $w_{l}$, where $w_{h}\geq w_{l}>0$.
Lines 4-14 describe an iterative random walk to select nodes until $B$ is reached. During each iteration, the algorithm chooses a node $v$ from $L$ based on the probability distribution $L_{w}$ (Line 6). Next, it picks a neighbor $u$ of $v$ using a weighted selection process (Lines 7-9) (O1). Specifically, in Line 8, we steer the random walks towards relevant nodes, edges, or paths. The transition probability matrices for node, edge, and path hypotheses are shown in Figures  3a,  3b, and  3c, respectively. For a node hypothesis, a higher weight $w_{h}$ is assigned to $x_{1}$ regardless of the current node. For an edge hypothesis, $w_{h}$ is given to $x_{2}$ when the current node is $x_{1}$, and given to $x_{1}$ in other cases. This increases the chance of sampling more relevant edges $(x_{1},x_{2})$. The random walks are 1st-order in Figures  3a and  3b. Path hypotheses ($l=2$) have 2nd-order random walks (Figure  3c), meaning the transition depends on the current and previous nodes. For example, given two nodes $x_{1}$ and $x_{2}$, $x_{3}$ gets a higher weight $w_{h}$. The consistent $w_{h}$ prevents sampling bias within the population. After adding nodes $u$ and $v$ to $\mathcal{V}_{S}$ (Line 10), the algorithm updates $L$ by replacing $v$ with $u$ (Lines 11) and decreases the sampling budget by one (Lines 12). This iteration continues until $B>m$. The resulting sampled graph $\mathcal{S}$ is the induced subgraph from $\mathcal{V}_{S}$.
$x_{1}$$y$$x_{1}$$w_{h}$$w_{l}$$y$$w_{h}$$w_{l}$(a)$x_{1}$$x_{2}$$y$$x_{1}$$w_{l}$$w_{h}$$w_{l}$$x_{2}$$w_{h}$$w_{l}$$w_{l}$$y$$w_{h}$$w_{l}$$w_{l}$(b)$x_{1}$$x_{2}$$x_{3}$$y$$x_{1},x_{1}$$w_{l}$$w_{h}$$w_{l}$$w_{l}$$x_{1},x_{2}$$w_{l}$$w_{l}$$w_{h}$$w_{l}$$x_{1},x_{3}$$w_{h}$$w_{l}$$w_{l}$$w_{l}$$x_{1},y$$w_{h}$$w_{l}$$w_{l}$$w_{l}$(c)
Figure 3. Transition probability matrices ${Q}$ for (a) node, (b) edge, and (c) path $(l=2)$ hypotheses (up to the first four rows). $x_{i}$ represents nodes in $\mathcal{G}$ satisfying the $i$-th node modifier on $\mathcal{P}$ and $y$ represents other nodes.